OpenStudy (anonymous):

The legs of an isosceles right triangle each measure 10 inches. What is the length of the hypotenuse of this triangle, to the nearest tenth of an inch? a. 6.3 b. 7.1 c. 14.1 4. 17.1 Can someone please go over this with me? And if there is a formula to solve this problem, please tell me it. Thank you!

OpenStudy (anonymous):

let the hyp be h and leg l we know i a right isos H=root2*l soo

OpenStudy (mrnood):

 You should know the 'Pythagoras formula' to allow you to calculate H

OpenStudy (anonymous):

root2 * 10 will give 14. something which is C :)

OpenStudy (mrnood):

However, this is a special case of a right angle triangle - being isosceles so whatever the side we have H^2 = 2 L^2 or H = sqrt (2L^2) therefore H =L* sqrt 2